# Experiment design for continuous-time systems using generalized filtering

### v1

@@ Abstract
The goal of experiment design is to select the inputs of a dynamical system in such a way that the resulting data contain sufficient information for data-driven modeling and control. This paper investigates the problem of experiment design for continuous-time systems under piecewise constant input signals. To obviate the need for measuring time derivatives of (data) trajectories, we introduce a generalized filtering framework. Our main result is to establish conditions on the input and the filter functions under which the filtered data are informative for system identification, i.e., they satisfy a certain rank condition. The proposed framework encompasses several filter functions that have already appeared in the literature. Building on the proposed filtering framework, we develop an experiment design procedure where the input signal is designed online during system operation. This method is shown to be sample efficient, in the sense that it deals with the least possible number of filtered data samples for system identification.
@@
thanks: This paper was not presented at any IFAC meeting. Corresponding author: Simone Baldi. This work was partially supported by Jiangsu Provincial Scientific Research Center of Applied Mathematics grant BK20233002. H. J. van Waarde acknowledges financial support by the Dutch Research Council under the NWO Talent Programme Veni Agreement (VI.Veni.222.335).
@@ 1 Introduction
The growing complexity of modern engineering systems complicates modeling from first principles and motivates the use of data for system modeling and control [de2019formulas, van2023informativity, DBLSCT2025]. Data-driven techniques utilize measurements collected from the system to model the system itself or to design controllers. Data-driven techniques for discrete-time systems [de2019formulas, van2023informativity, DBLSCT2025, van2020data, berberich2021data, rotulo2022online, wang2025necessary, eising2025data, zhao2025data] have a richer development as compared to their continuous-time counterparts [berberich2021datacontinuous, rapisarda2023orthogonal, ohta2024data, eising2025sampling], partly due to the challenge of accurately measuring time derivatives of trajectories in continuous-time systems. One of the possible approaches to address this challenge is to apply one of the various filtering methods proposed in the literature [ohsumi2002subspace, ohta2011stochastic, ohta2024data]. Alternative methods involve, for instance, robust adaptation [yucelen2011derivative], estimation using discretization [berberich2021datacontinuous] and orthogonal polynomial bases [rapisarda2023orthogonal].
While most of the existing literature performs analysis and control using given datasets, other studies address a preliminary question—the design of input signals, with the aim of generating suitable data. This phase, known as experiment design, is crucial to obtain data containing sufficient information for data-driven modeling and control. The theoretical foundations for experiment design are provided by the fundamental lemma by Willems et al., originally formulated for discrete-time systems [willems2005note]. The lemma characterizes those input signals resulting in collected data that represent all possible trajectories of the system. Different extensions of this lemma to continuous-time systems have been proposed [rapisarda2022persistency, lopez2022continuous, rapisarda2023fundamental, lopez2024input], demonstrating from different perspectives that a single persistently exciting input signal can generate a system trajectory rich enough to reproduce all other trajectories. Nevertheless, these results typically rely on direct access to a number of time derivatives of (data) trajectories. The experiment design for continuous-time systems without the availability of time derivatives remains a challenge.
This naturally raises the question of whether such a problem can be addressed by leveraging on filtering methods like those in [ohsumi2002subspace, ohta2011stochastic, ohta2024data]. This work gives a positive answer to this question by means of a generalized filtering framework encompassing existing filtering methods. Answering this question is far from trivial, as no existing result in the literature guarantees that the data obtained after filtering will retain sufficient information for data-driven modeling and control. In this work, we provide conditions under which the filtered data are informative for system identification, i.e., they satisfy a certain rank condition.
@@
In this work, we introduce a generalized filtering framework (cf. Eq. (1)) which encompasses several filtering methods from the literature, see Example 1. Based on this framework, the two main contributions of the paper are as follows:
  • We establish a relation between the filtered data and the sampled data (cf. Eq. (32)). This relation enables us to establish Proposition 1, which provides conditions under which the filtered data are informative for system identification. Notably, the conditions we present are missing in all filtering methods in the literature, which rely on assuming informativity of the filtered data, rather than deriving the conditions making informativity possible.
  • We propose an online experiment design method (cf. Proposition 3) that guarantees the filtered data to be informative with the minimum possible number of samples. Theorem 2 shows that the designed input signal is such that the data captures the system’s dynamics at all times between sampling instants, which establishes a connection between the proposed method and the continuous-time Willems et al.’s fundamental lemma in [lopez2022continuous].
@@
1.1 Notation
We denote the set of non-negative integers by $\mathbb{Z}_{+}$, the set of positive integers by $\mathbb{N}$, and the set of non-negative real numbers by $\mathbb{R}_{+}$. Given a matrix $A\in\mathbb{R}^{n\times m}$, its Moore-Penrose pseudo-inverse is denoted by $A^{\dagger}$. The identity matrix with appropriate dimensions is denoted by $I$. Given vectors $v_{k},v_{k+1},\dots,v_{\ell}\in\mathbb{R}^{n}$ with $k\leq\ell$, we define $v_{[k,\ell]}:=[v_{k}\ v_{k+1}\ \cdots\ v_{\ell}]\in\mathbb{R}^{n\times(\ell-k+1)}$.
@@ 2 Generalized filtering and filtered data
For $\mathcal{T}>0$ and $\ell\in\{1,2,\dots,M\}$, we introduce the filter functions $g_{\ell}:[0,\mathcal{T})\to\mathbb{R}$. We let $0=t_{0}<t_{1}<t_{2}<\cdots<t_{q}=\mathcal{T}$ and assume that $g_{\ell}$ is continuously differentiable on $t_{j-1},t_{j})$ for $j=1,2,\dots,q$. Thus, $g_{\ell}$ is a piecewise continuously differentiable function. In addition, we will assume that $g_{\ell}(t_{j}^{-}):=\lim_{t\uparrow t_{j}}g_{\ell}(t)$ exists for all $j=1,2,\dots,q$. The functions $g_{\ell}$ are used to produce filtered data $w^{\rm f}_{\ell}$ from a given signal $w:[0,\mathcal{T})\to\mathbb{R}^{n}$. More precisely, we define
@@
Note that (1) can accommodate several filtering methods from the literature as special cases. For instance, in [ohta2024data], the following set of filtered data was considered:
@@
and $h_{j}:[0,\mathcal{T})\to\mathbb{R}$ and $s_{k}\in[0,\mathcal{T})$ for $j=1,2,\dots,p$ and $k=1,2,\dots,r$. Note that $\mathcal{D}$ can be captured by (1) by setting $M=pq$ and defining
@@
for $j\in\{1,\dots,p\}$ and $k\in\{1,\dots,q\}$. The generality of (1) is further elaborated in the next example.
@@
    compactly supported test functions [ohsumi2002subspace] such as:
@@
    first Laguerre basis function [ohta2024data]:
@@
    low-pass filter function [roy2017combined]:
@@
(b) $g_{\ell}(t)=e^{-\frac{1}{1-(t-\ell+1)^{2}}}$
(c) $g_{\ell}(t)=\sqrt{2}e^{-(t-\ell+1)}$
(d) $g_{\ell}(t)=e^{t-\ell}$
Figure 1: Examples of filter functions for $\ell\in\{1,2,3\}$
Fig. 1(a)-1(d) show the four filter functions in (2)-(5) with parameters $\rho=1$, $T=1$ and $\ell\in\{1,2,3\}$. In some cases, the filtered data in (1) can be interpreted as the state of a dynamical system sampled at certain time instants. To illustrate this, consider the low-pass filter function (5). For any signal $w:[0,\mathcal{T})\to\mathbb{R}^{n}$, low-pass filtering can be described by the differential equation
@@
where $w^{\rm f}(t)\in\mathbb{R}^{n}$ denotes the filtered signal. The solution of (6) is
@@
where we have used (1), and the specific low pass filtering function $g_{\ell}$ defined in (5). We conclude that the filtered data $w_{\ell}^{f}$ can be interpreted as the state of the dynamical system (6) at specific time instants. The low-pass filter function (5) is often employed to avoid the need for measuring the time derivative of (state) variables [roy2017combined, cho2017composite]. Indeed, denote
@@ 3 Problem formulation
where $x(t)\in\mathbb{R}^{n}$ is the system state, $u(t)\in\mathbb{R}^{m}$ is the control input, and $A\in\mathbb{R}^{n\times n},B\in\mathbb{R}^{n\times m}$ are the system matrices, assumed to be unknown. Throughout the paper, we do assume that the pair $(A,B)$ is controllable.
Let $\mathcal{AC}([0,\mathcal{T}),\mathbb{R}^{n})$ be the space of absolutely continuous functions from $0,\mathcal{T})$ to $\mathbb{R}^{n}$, and let $\mathcal{PC}([0,\mathcal{T}),\mathbb{R}^{m})$ be the space of piecewise continuous functions from $0,\mathcal{T})$ to $\mathbb{R}^{m}$. We recall that $x\in\mathcal{AC}([0,\mathcal{T}),\mathbb{R}^{n})$ is a Carathéodory solution of (10) on $0,\mathcal{T})$ if (10a) holds for almost all $t\in[0,\mathcal{T})$ and (10b) is satisfied. We define the behavior [willems1997introduction] of (10) as
@@
where $x(t_{j}^{-})$ is well defined for all $j\in\{1,2,\dots,q\}$ because $x$ is absolutely continuous. This shows that there is no need for measuring the time derivative $\dot{x}$ of the state in order to obtain the filtered time derivative $x^{\rm df}_{\ell}$. Indeed, by (15), $x_{\ell}^{\rm df}$ may be obtained directly from measurements of $x$ and the filter function $g_{\ell}$. Since integration is a linear transformation, (12)-(14) and (10) yield the algebraic equation
@@
The rank condition (17) is crucial for data-driven modeling and control based on filtered data, see e.g., [ohta2024data]. Indeed, as an example we note that (16) and (17) imply that the system matrices $A$ and $B$ can be uniquely recovered from the data as
@@
i.e., the filtered data $(x^{\rm df}_{[1,M]},x^{\rm f}_{[1,M]},u^{\rm f}_{[1,M]})$ are informative for system identification. Analogous versions of the rank condition in (17) have also appeared for other types of data in [de2019formulas, berberich2021datacontinuous, rapisarda2023orthogonal], for example, for samples of continuous-time trajectories [de2019formulas, berberich2021datacontinuous] and coefficient matrices corresponding to certain basis functions [rapisarda2023orthogonal].
@@ 4 Rank condition for filtered data
Let $T>0$ and $N\in\mathbb{N}$ and define $\mathcal{T}:=NT$. We consider a piecewise constant input $u:[0,\mathcal{T})\to\mathbb{R}^{m}$, i.e.,
@@ 4.1 Continuous-time fundamental lemma
For a piecewise constant input, exact discretization of the continuous-time system (10) with sampling time $T$ results in the discrete-time dynamics [chen2012optimal]
@@
In order to ensure that the discrete-time system (22) preserves the controllability property of the original continuous-time system (10), the following assumption from the literature is considered.
@@
Let $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$ be the eigenvalues of $A$. Then, for any distinct $j,l\in\{1,2,\dots,n\}$ and $q\in\mathbb{N}$, the sampling time $T$ is such that
@@
As highlighted in [chen2012optimal], Assumption 1 is mild in the sense that (24) holds for all $T$ except for a set of measure zero. Based on Assumption 1, we recall a lemma asserting that controllability is preserved under discretization.
@@
([chen2012optimal, Theorem 3.2.1]). Let Assumption 1 hold. Then $(A_{T},B_{T})$ is controllable if $(A,B)$ is controllable.
Using this property, we recall the continuous-time fundamental lemma from [lopez2022continuous].
@@
([lopez2022continuous, Lemma 1]). Let Assumption 1 hold. Consider the trajectory $(u,x)\in\mathfrak{B}_{NT}$ with $u$ a piecewise constant input signal in (19) satisfying $\operatorname{rank}(\mathcal{H}_{n+1}(\mu_{[0,N-1]}))=(n+1)m,$, where
@@
Lemma 2 provides an experiment design method to achieve (25) for the sampled data (20)-(21). Yet, we stress on the fact that nothing can be concluded about the rank condition (17) for the filtered data, as required to solve Problem 1 and 2. Hence, we proceed by analyzing the relation between the sampled and the filtered data under piecewise constant input signals.
@@ 4.2 Rank relation between sampled and filtered data
    $g(t)=\rho t^{2}(T-t)^{2}\text{ or }g(t)=e^{-\frac{bT^{2}}{T^{2}-t^{2}}},$ (27)
@@
Proof. We first show that $\bar{A}$ is nonsingular. We consider the Jordan normal form of $A$, i.e., $A=P\Lambda P^{-1}$, where $P\in\mathbb{R}^{n\times n}$ is nonsingular,
$\Lambda=\!\begin{bmatrix}\Lambda_{1}&0&\cdots&0\\ 0&\Lambda_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\Lambda_{k}\end{bmatrix}\!,\ \Lambda_{l}=\!\begin{bmatrix}\lambda_{l}&1&\cdots&0&0\\ 0&\lambda_{l}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\lambda_{l}&1\\ 0&0&\cdots&0&\lambda_{l}\\ \end{bmatrix}\!\in\mathbb{R}^{n_{l}\times n_{l}},$
$l\in\{1,2,\dots,q\}$ and $\sum_{l=1}^{q}n_{l}=n$. Then,
@@
Note that for any ${l}\in\{1,2,\dots,q\}$,
@@
([bernstein2009matrix, Fact 2.10.14]) For any $P\in\mathbb{R}^{p\times q}$ and $Q\in\mathbb{R}^{q\times r}$,
@@
The latter condition is, in general, difficult to impose by choosing the inputs $\mu_{0},\mu_{1},....,\mu_{N-1}$. Indeed, the reason is that the states of the system cannot be chosen freely, but depend on the choice of inputs and the (unknown) dynamics (10). Therefore, it is in general not possible to design an experiment so that the kernel of $\bar{D}$ coincides with a given subspace. We do note, however, that the problem simplifies in the case that $N=n+m$. In this case, if $\operatorname{rank}(\bar{D})=\operatorname{rank}(\bar{F})=n+m$, then $\ker(\bar{D})=\{0\}$ and thus $\operatorname{rank}(\bar{D}\bar{F})=n+m$. This suggests that it is important to design experiments of length precisely $N=n+m$, which will be explored in Section 5. In what follows, we summarize our progress in a proposition.
@@
Let Assumption 2 hold. If $N=n+m$, $\operatorname{rank}(\bar{F})=n+m$ and
$\operatorname{rank}\left(\begin{bmatrix}\chi_{[0,N-1]}\\ \mu_{[0,N-1]}\end{bmatrix}\right)=n+m,$ (36)
then (17) holds.
@@
Consider a trajectory $(u,x)\in\mathfrak{B}_{NT}$. Let the functions $g_{\ell}$, for $\ell=1,2,\dots,M$, satisfy Assumption 2 with $f_{\ell}((\ell-1)T)\neq 0$ and $f_{\ell}(kT)=0$ for all $k\geq\ell$ and $\ell\in\{1,2,\dots,\min\{N,M\}\}$. Then,
@@
In the next section, we will focus on an experiment design method that imposes the rank condition (36) by appropriate choice of the inputs. This, in combination with Proposition 2, will lead to the desired rank condition on the filtered data as in (17).

### v2

# Online experiment design for continuous-time systems using generalized filtering

@@ Abstract
The goal of experiment design is to select the inputs of a dynamical system in such a way that the resulting data contain sufficient information for system identification and data-driven control. This paper investigates the problem of experiment design for continuous-time systems under piecewise constant input signals. To obviate the need for measuring time derivatives of (data) trajectories, we introduce a generalized filtering framework. Our main result is to establish conditions on the input and the filter functions under which the filtered data are informative for system identification, i.e., they satisfy a certain rank condition. We assume that the filter functions are piecewise continuously differentiable, encompassing several filter functions that have appeared in the literature. Building on the proposed filtering framework, we develop an experiment design procedure, adapted from experiment design results for discrete-time systems, where the piecewise constant input signal is designed online during system operation. This method is shown to be sample efficient, in the sense that it deals with the least possible number of filtered data samples for system identification.
@@
thanks: This paper was not presented at any IFAC meeting. This work was partially supported by Jiangsu Provincial Scientific Research Center of Applied Mathematics grant BK20233002. H. J. van Waarde acknowledges financial support by the Dutch Research Council under the NWO Talent Programme Veni Agreement (VI.Veni.222.335). Corresponding author: Simone Baldi.
@@ 1 Introduction
The growing complexity of modern engineering systems complicates modeling from first principles and motivates the use of data for modeling and control [6, 21, 20]. Data-driven techniques utilize measurements collected from the system to model the system itself or to design controllers. Data-driven techniques for discrete-time systems [6, 21, 20, 22, 1, 18, 24, 8, 28] have a richer development as compared to their continuous-time counterparts [2, 17, 14, 7], partly due to the challenge of accurately measuring time derivatives of trajectories in continuous-time systems. One of the possible approaches to address this challenge is to apply one of the various filtering methods proposed in the literature [12, 13, 14]. Alternative methods involve, for instance, robust adaptation [27], estimation using discretization [2] and orthogonal polynomial bases [17].
While most of the existing literature performs analysis and control using given datasets, other studies address a preliminary question—the design of input signals, with the aim of generating suitable data. This phase, known as experiment design, is crucial to obtain data containing sufficient information for system identification and data-driven control. The theoretical foundations for experiment design are provided by the fundamental lemma by Willems et al., originally formulated for discrete-time systems [26]. The lemma characterizes those input signals resulting in collected data that represent all possible trajectories of the system. Different extensions of this lemma to continuous-time systems have been proposed [15, 10, 16, 9], demonstrating from different perspectives that a single persistently exciting input signal can generate a system trajectory rich enough to reproduce all other trajectories. Nevertheless, these results typically rely on direct access to a number of time derivatives of (data) trajectories. The experiment design for continuous-time systems without the availability of time derivatives remains a challenge.
This naturally raises the question of whether such a problem can be addressed by leveraging on filtering methods like those in [12, 13, 14]. This work gives a positive answer to this question by means of a generalized filtering framework encompassing existing filtering methods. Answering this question is far from trivial, as no existing result in the literature guarantees that the data obtained after filtering will retain sufficient information for system identification and data-driven control. In this work, we provide conditions under which the data obtained after filtering are informative for system identification, i.e., they satisfy a certain rank condition that guarantees unique identifiability of the system.
@@
In this work, we introduce a generalized filtering framework (see Eq. (1)) which encompasses several filtering methods from the literature, see Example 1. Based on this framework, the two main contributions of the paper are as follows:
  • We establish a relation between the filtered data and the sampled data (see Eq. (32)). This relation enables us to establish Proposition 1, which provides conditions under which the filtered data are informative for system identification. Notably, the conditions we present are missing in all filtering methods in the literature, which rely on assuming informativity of the filtered data, rather than deriving the conditions making informativity possible.
  • Based on the discrete-time setting in [23] we develop an online experiment design method for continuous-time systems (see Proposition 3) and we show that such method guarantees the filtered data to be informative with the minimum possible number of samples (see Theorem 1). Theorem 2 shows that the designed input signal is such that the data captures the system’s dynamics at all times between sampling instants, which establishes a connection between the proposed method and the continuous-time Willems et al.’s fundamental lemma in [10].
@@
Notation: We denote the set of non-negative integers by $\mathbb{Z}_{+}$, the set of positive integers by $\mathbb{N}$, and the set of non-negative real numbers by $\mathbb{R}_{+}$. Given a matrix $A\in\mathbb{R}^{n\times m}$, its Moore-Penrose pseudo-inverse is denoted by $A^{\dagger}$. The identity matrix with appropriate dimensions is denoted by II. Given vectors $v_{k},v_{k+1},\dots,v_{\ell}\in\mathbb{R}^{n}$ with $k\leq\ell$, we define $v_{[k,\ell]}:=[v_{k}\ v_{k+1}\ \cdots\ v_{\ell}]\in\mathbb{R}^{n\times(\ell-k+1)}$.
@@ 2 Generalized filtering and filtered data
For $\mathcal{T}>0$, we introduce $M$ filter functions $g_{\ell}\!:[0,\mathcal{T})\!\to\!\mathbb{R}$ for $\ell\in\{1,2,\dots,M\}$. We let $0=t_{0}<t_{1}<t_{2}<\cdots<t_{q}=\mathcal{T}$ and assume that $g_{\ell}$ is continuously differentiable on $t_{j-1},t_{j})$ for $j=1,2,\dots,q$. Thus, $g_{\ell}$ is a piecewise continuously differentiable function. In addition, we assume that $g_{\ell}(t_{j}^{-}):=\lim_{t\uparrow t_{j}}g_{\ell}(t)$ exists for all $j=1,2,\dots,q$. The functions $g_{\ell}$ are used to produce filtered data $w^{\rm f}_{\ell}$ from a given signal $w:[0,\mathcal{T})\to\mathbb{R}^{n}$. More precisely, we define
@@
Remark 1.
Note that (1) encompasses several filtering methods from the literature. For instance, in [14], the following set of filtered data was considered:
@@
and $h_{j}:[0,\mathcal{T})\to\mathbb{R}$ and $s_{k}\in[0,\mathcal{T})$ for $j=1,2,\dots,p$ and $k=1,2,\dots,r$. Indeed, (1) can accommodate $\mathcal{D}$ by setting $M=pr$ and defining
@@
for $j\in\{1,2,\dots,p\}$ and $k\in\{1,2,\dots,r\}$.
The generality of (1) is further elaborated in the next example.
@@
    compactly supported test functions [12] such as:
@@
    first Laguerre basis function [14]:
@@
    low-pass filter function [19]:
@@
(b) $g_{\ell}(t)=e^{\text{\textminus}\frac{1}{1\text{\textminus}(t\text{\textminus}\ell+1)^{2}}}$
(c) $g_{\ell}(t)=\sqrt{2}e^{\ell\text{\textminus}1\text{\textminus}t}$
(d) $g_{\ell}(t)=e^{t\text{\textminus}\ell}$
Figure 1: Examples of filter functions for $\ell\in\{1,2,3\}$: (a)-(b) compactly supported test functions, (c) first Laguerre basis function, (d) low-pass filter function.
Figures 1(a)-1(d) show the four filter functions in (2)-(5) with parameters $\rho=1$, $T=1$ and $\ell\in\{1,2,3\}$. In the low-pass filtering case, the filtered data in (1) can be interpreted as the state of a dynamical system sampled at certain time instants. To illustrate this, consider the dynamical system
@@
where $w^{\rm f}(t)\in\mathbb{R}^{n}$ is the state of system (6). The solution of (6) is
@@
where we have used (1), and the specific low-pass filtering function $g_{\ell}$ in (5). We conclude that the filtered data $w_{\ell}^{f}$ obtained from the low-pass filter function (5) are samples of the state of the dynamical system (6) at specific time instants. The low-pass filter function (5) is often employed to avoid the need for measuring the time derivative of (state) variables [19, 5]. Indeed, denote
@@ 3 Problem formulation
where $x(t)\in\mathbb{R}^{n}$ is the system state, $u(t)\in\mathbb{R}^{m}$ is the control input, and $A\in\mathbb{R}^{n\times n},B\in\mathbb{R}^{n\times m}$ are the system matrices, assumed to be unknown. Throughout the paper, we assume that the pair $(A,B)$ is controllable.
Let $\mathcal{AC}([0,\mathcal{T}),\mathbb{R}^{n})$ be the space of absolutely continuous functions from $0,\mathcal{T})$ to $\mathbb{R}^{n}$, and let $\mathcal{PC}([0,\mathcal{T}),\mathbb{R}^{m})$ be the space of piecewise continuous functions from $0,\mathcal{T})$ to $\mathbb{R}^{m}$. We recall that $x\in\mathcal{AC}([0,\mathcal{T}),\mathbb{R}^{n})$ is a Carathéodory solution of (10) on $0,\mathcal{T})$ if (10a) holds for almost all $t\in[0,\mathcal{T})$ and (10b) is satisfied. We define the behavior [25] of (10) as
@@
where $x(t_{j}^{-})$ is well defined for all $j\in\{1,2,\dots,q\}$ because $x$ is absolutely continuous. This shows that there is no need for measuring the time derivative $\dot{x}$ of the state in order to obtain the filtered time derivative $x^{\rm df}_{\ell}$.
Since integration is a linear transformation, (12)-(14) and (10) yield the algebraic equation
@@
The rank condition (17) is crucial for system identification and data-driven control based on filtered data, see e.g., [14]. Indeed, as an example we note that (16) and (17) imply that the system matrices AA and $B$ can be uniquely recovered from the data as
@@
i.e., the filtered data $(x^{\rm df}_{[1,M]},x^{\rm f}_{[1,M]},u^{\rm f}_{[1,M]})$ are informative for system identification. Analogous versions of the rank condition in (17) have also appeared for other types of data in [6, 2, 17], for example, for samples of continuous-time trajectories [6, 2] and coefficient matrices corresponding to certain basis functions [17].
@@ 4 Rank condition for filtered data
Let $\mathcal{T}:=NT$ with $N\in\mathbb{N}$ and sampling time $T>0$. We consider a piecewise constant input trajectory $u:[0,\mathcal{T})\to\mathbb{R}^{m}$, defined by
@@ 4.1 Continuous-time fundamental lemma
For a piecewise constant input, exact discretization of the continuous-time system (10) with sampling time $T$ results in the discrete-time dynamics [4]
@@
To ensure that the discrete-time system (22) preserves the controllability property of the original continuous-time system (10), the following assumption from the literature is considered.
@@
Let $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$ be the eigenvalues of AA. Then, for any distinct $j,l\in\{1,2,\dots,n\}$ and $q\in\mathbb{N}$, the sampling time $T$ is such that
@@
As highlighted in [4], Assumption 1 is mild in the sense that (24) holds for all $T$ except for a set of measure zero. Based on Assumption 1, we recall a lemma asserting that controllability is preserved under discretization.
@@
([4, Theorem 3.2.1]). Let Assumption 1 hold. Then $(A_{T},B_{T})$ is controllable if $(A,B)$ is controllable.
Using this property, we recall the continuous-time fundamental lemma from [10].
@@
([10, Lemma 1]). Let Assumption 1 hold. Consider the trajectory $(u,x)\in\mathfrak{B}_{NT}$ with $u$ a piecewise constant input signal in (19) satisfying $\operatorname{rank}(\mathcal{H}_{n+1}(\mu_{[0,N-1]}))=(n+1)m$, where
@@
Lemma 2 provides an experiment design method to achieve (25) for the sampled data (20)-(21). Yet, we stress on the fact that nothing can be concluded about the rank condition (17) for the filtered data, as required to solve Problems 1 and 2. Hence, we proceed by analyzing the relation between the sampled and the filtered data under piecewise constant input signals.
@@ 4.2 Rank relation between sampled and filtered data
    $g(t)=\rho t^{2}(T-t)^{2}\text{ or }g(t)=e^{-\frac{bT^{2}}{T^{2}\text{\textminus}t^{2}}},$ (27)
@@
Proof. We first show that $\bar{A}$ is nonsingular. We consider the Jordan normal form of AA, i.e., $A=P\Lambda P^{-1}$, where $P\in\mathbb{R}^{n\times n}$ is nonsingular,
$\Lambda=\!\begin{bmatrix}\Lambda_{1}&0&\cdots&0\\ 0&\Lambda_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\Lambda_{p}\end{bmatrix}\!,\ \Lambda_{l}=\!\begin{bmatrix}\lambda_{l}&1&\cdots&0&0\\ 0&\lambda_{l}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\lambda_{l}&1\\ 0&0&\cdots&0&\lambda_{l}\\ \end{bmatrix}\!\in\mathbb{R}^{n_{l}\times n_{l}},$
$p$ is the number of Jordan blocks, $l\in\{1,2,\dots,p\}$ and $\sum_{l=1}^{p}n_{l}=n$. Then,
@@
Note that for any ${l}\in\{1,2,\dots,p\}$,
@@
([3, Fact 2.10.14]) For any $P\in\mathbb{R}^{n_{1}\times n_{2}}$ and $Q\in\mathbb{R}^{n_{2}\times n_{3}}$,
@@
The latter condition is, in general, difficult to impose by choosing the inputs $\mu_{0},\mu_{1},....,\mu_{N-1}$. The reason is that the states of the system cannot be chosen freely, but depend on the choice of inputs and the (unknown) dynamics (10). Therefore, it is in general not possible to design an experiment so that the kernel of $\bar{D}$ coincides with a given subspace. We do note, however, that the problem simplifies in the case that $N=n+m$. In this case, if $\operatorname{rank}(\bar{D})=\operatorname{rank}(\bar{F})=n+m$, then $\ker(\bar{D})=\{0\}$ and thus $\operatorname{rank}(\bar{D}\bar{F})=n+m$. In what follows, we summarize our progress in a proposition, which provides a solution to Problem 1.
@@
Let Assumption 2 hold. Suppose that $N=n+m$ and $\operatorname{rank}(\bar{F})=n+m$. Then, (17) holds if and only if $\operatorname{rank}(\bar{D})=n+m$.
Proof. By (32), the necessity holds since
$\operatorname{rank}\left(\begin{bmatrix}x_{[1,M]}^{f}\\ u_{[1,M]}\end{bmatrix}\right)\leq\operatorname{rank}(\bar{D})\leq n+m.$
For sufficiency, $N=n+m$ implies $\ker(\bar{D})=\{0\}$. Then, by Lemma 4, $\operatorname{rank}(\bar{D}\bar{F})=\operatorname{rank}(\bar{F})=n+m$. Since $\bar{C}\in\mathbb{R}^{(n+m)\times(n+m)}$, Lemma 3 implies that (17) holds. $\Box$
@@
Let Assumption 2 hold. Suppose that the functions $g_{\ell}$, for $\ell=1,2,\dots,M$, satisfy $f_{\ell}((\ell-1)T)\neq 0$ and $f_{\ell}(kT)=0$ for all $k\geq\ell$ and $\ell\in\{1,2,\dots,\min\{N,M\}\}$. Then,
@@
Remark 2.
Note that the condition $f_{\ell}(kT)=0$ for all $k\geq\ell$ in Proposition 2 implies that the signal $x$ (or $u$) over the interval $\ell T,\mathcal{T})$ does not affect the filtered data $x^{\rm f}_{\ell}$ (or $u^{\rm f}_{\ell}$). By examining the matrix $\bar{F}$ in Example 3 for the filter functions (27)-(29), it can be observed that the result of Proposition 2 applies to (27) and (29), but not to (28).
In the next section, we will focus on an experiment design method that imposes $\operatorname{rank}(\bar{D})=n+m$ by appropriate choice of the inputs. This, in combination with Proposition 2, will lead to the desired rank condition on the filtered data as in (17).